A company is investigating the feasibility of purchasing a machine at the present time for $200,000. The following information pertains to the acquisition and use of the machine over the next five years:

In year one the cash revenues generated by the machine are expected to be $76,000, and these revenues are expected to increase by 6% year-over-year in each of the remaining 4 years that they own the machine.
The machine is expected to be sold at the end of 5 years for its estimated salvage value of $20,000.
Cash operating expenses directly associated with the generation of the annual revenues noted above are expected to be 65% of the revenues in year one, 64% of the revenues in year two, 63% of the revenues in year three, 62% of the revenues in year four, and 61% of the revenues in year five. In addition to the cash operating expenses, there will also be an annual depreciation expense of one-fifth of the cost (minus the salvage value) of the asset each year for the 5 years that the asset is held.
The machine will require a major overhaul costing $40,000 at the end of year three.

Prepare a cash flow spreadsheet to summarize the information above. For simplicity purposes you can ignore income taxes in this problem.
Answer #1
Cash flow details of the feasibility of purchase of machinery
Year Cash revenue Cash operating expenses (outflow) Cash flow (outflow)/inflow related to asset Total cash inflow/(outflow)
0                                                         (200,000)                               (200,000)
1                                                                 76,000                                                           (49,400)                                  26,600
2                                                                 80,560                                                           (51,558)                                  29,002
3                                                                 85,394                                                           (53,798)                                                          (40,000)                                  (8,404)
4                                                                 90,517                                                           (56,121)                                  34,397
5                                                                 95,948                                                           (58,528)                                                            20,000                                  57,420
Details of cash revenue generated by the Machine
Year Cash revenue Details
1                                                                 76,000 First year cash revenue
2                                                                 80,560 Increase by 6% every year ( 76,000*106%)
3                                                                 85,394 Increase by 6% every year ( 80,560*106%)
4                                                                 90,517 Increase by 6% every year ( 85,394*106%)
5                                                                 95,948 Increase by 6% every year ( 790,517*106%)
Details of cash operating expenses
Year cash operating expenses Details Computation
1                                                                 49,400 65% of revenue in the first year =76,000*65%
2                                                                 51,558 64% of revenue in the second year =80,560*64%
3                                                                 53,798 63% of revenue in the third year =85,394*63%
4                                                                 56,121 62% of revenue in the fourth year =90,517*62%
5                                                                 58,528 61% of revenue in the fifth year =95,948*61%
Details of cash floe related to machine
Year Cash flow (outflow)/inflow related to asset
0                                                              (200,000)
1
2
3                                                               (40,000)
4
5                                                                 20,000
Details of depreciation
Since in the question it has been mention to ignore the income tax therefore there will not be any impact for depreciation as we are not considering the tax benefit on depreciation and it is assumed that over hauling expenses are revenue in nature and will be charge off to profit and loss account
Year Depreciation on originally purchased asset
=Cost of machine-Salvage value/life of the asset)
=200,000-20,000)/5
1                                                                 36,000
2                                                                 36,000
3                                                                 36,000
4                                                                 36,000
5                                                                 36,000